OpenStudy (alexh107):

Simplify (sqrt5 -2i)^2

OpenStudy (danieldjpon3):

uh

OpenStudy (raffle_snaffle):

What are we solving? For i? Can we see a screen shot of the problem?

OpenStudy (alexh107):

\[(\sqrt{5}-2i)^2\]

OpenStudy (alexh107):

I should have said simplify I think, my mistake.

OpenStudy (sshayer):

\[\left( \sqrt{5}-2 \iota \right)^2=5+4 \iota ^2-4\sqrt{5} \iota =5-4-4\sqrt{5}\iota =1-4 \sqrt{5} \iota \]

OpenStudy (alexh107):

Ah okay, I'll look over that. Thank you!